• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name perdrisa

    Log entry time 16:00:29 on January26,2002

    Entry number 80886

    keyword=day shift log

    No beam from start of day shift, 8:00, to 9:00
    continue on kin. 4H, run 1700, started at 9:10. current 16 mA, thresh. 550 MeV
    from run 1702, start. at 9:44, current 10 mA.
    from run 1705, start at 11:20, current 19 mA, thresh. 650 MeV,
    run 1706 no data
    11:35, beam to MCC for tuning (prepare for larger currents); completed at 11:52
    12:00 MCC test to 50 mA ok
    12:00 to ....TS2 problems
    calo HV problems; 1 crate burned out (13:10); to be replaced (Eugene).
    HV problems fixed at 14:30
    usable beam back in hall at 14:45
    start run 1710 at 14:45.
    ROC 14 failure during 1711
    15:45, data stream stops, no immediate explanation
    15:55 ROC 4 failure